### v1

# Revisiting institutional punishment in the $N$-person prisoner’s dilemma

Abstract
The conflict between individual and collective interests makes fostering cooperation in human societies a challenging task, requiring drastic measures such as the establishment of sanctioning institutions. These institutions are costly because they have to be maintained regardless of the presence or absence of offenders. Here, we propose realistic improvements to the standard $N$-person prisoner’s dilemma formulation with institutional punishment by eliminating overpunishment, requiring a minimum number of contributors to establish the sanctioning institution, and sharing the cost among them once this minimum number is reached. In addition, we focus on large groups or communities for which sanctioning institutions are ubiquitous. Using the replicator equation framework for an infinite population, we find that by sufficiently fining players who fail to contribute either to the public good or to the sanctioning institution, a population of contributors immune to invasion by these free riders can be established, provided that the contributors are sufficiently numerous. In a finite population, we use finite-size scaling to show that, for some parameter settings, demographic noise helps to fixate the strategy that contributes to the public good but not to the sanctioning institution even for infinitely large populations when, somewhat counterintuitively, its proportion in the initial population vanishes with a small power of the population size.
I Introduction
Understanding and promoting cooperation in human societies has been declared one of the greatest challenges for science in the twenty-first century Kennedy_2005 . The main issue is the conflict between individual and collective interests: either individuals contribute to public goods at a personal cost, or they free ride on the efforts of others Hardin_1968 ; Axelrod_1984 . This conflict between different levels of selection, which generally arises when considering social traits, was already noted by Darwin in his discussion of courage and self-sacrifice in humans, since these are traits that are counter-selected within a social group, but increase the group’s chances of survival in the case of inter-group fighting Darwin_1871 . This type of social dilemma has been extensively studied within the framework of evolutionary game theory Maynard_1982 .
The usual evolutionary game theoretical framework for studying cooperation is the $N$-person prisoner’s dilemma Fox_1978 . In a play group of $N>1$ individuals, each individual can decide whether to contribute a fixed amount $c>0$ to the common pool (public goods). This amount is multiplied by a factor of $r>1$ and then divided among the $N-1$ other players. Following the terminology of the $2$-person prisoner’s dilemma, we refer to individuals that contribute to the common pool as cooperators and those that do not as defectors. Clearly, in this scenario cooperators are doomed to extinction unless there is some positive assortment between them Hamilton_1975 , which is best illustrated by the green-beard effect Hamilton_1964 , or repeated interactions that allow the emergence of reputation and reciprocity Wang_2023 ; Quan_2024 ; Han_2024 ; Lai_2024 .
An obvious way to promote cooperation would be to punish defectors. In fact, this works well when it comes to eliminating defectors, but since punishment is costly, peer punishers (who also contribute to public goods) are also eliminated, leaving the resulting population of cooperators vulnerable to invasion by defectors Boyd_1992 . In this scenario, a peer punisher imposes a fixed penalty on each defector. Introducing second-order punishment, i.e., punishing those who do not punish, leads to the elimination of both defectors and cooperators. However, since second-order punishment is ineffective in the absence of defectors, cooperators may take over the population due to random drift, leaving it vulnerable to invasion by defectors again. One way out of this conundrum is to assume that the game is noncoercive, i.e., some players may abstain from participating, and to set the payoff of nonparticipants to an ad hoc value greater than the payoff received by a defector in a group of defectors, but smaller than the payoff received by a pair of cooperators Fowler_2005 ; Hauert_2007 . This guarantees that a population of defectors can be invaded by nonparticipants, which in turn can be invaded by cooperators, which can be invaded by defectors again, creating a sort of rock-paper-scissors game scenario Hannelore_2010 .
However, at least from the perspective of (modern) human social organization, a more natural way to deal with defectors of all kinds is to introduce pool or institutional punishment, where each punisher contributes a fixed amount in advance to a sanctioning institution Sigmund_2010 ; Szolnoki_2011 ; Perc_2012 . In particular, institutional punishers not only contribute $c$ to the public goods, but also advance an amount $\gamma$ to a sanctioning institution. Defectors are fined a fixed amount $\beta$ and cooperators are fined a smaller amount $\alpha\beta$ with $\alpha\in[0,1]$ for not contributing to the the sanctioning institution. This mimics the way contemporary societies are organized Ostrom_1990 . In particular, not participating in the game (e.g., not recognizing the right of the sanctioning institution to punish) is not an option.
Given the undeniably important role that institutional punishment plays in curbing defection in the real world, we consider here some realistic improvements to the standard $N$-person prisoner’s dilemma formulation with pool punishment Sigmund_2010 . In particular, we eliminate overpunishment (i.e., penalties for different types of offenses are fixed and do not depend on the wealth of the sanctioning institution Dercole_2013 ), but account for the high costs of creating and maintaining a sanctioning institution by requiring a minimum number of punishers to establish it. However, the maintenance costs are divided among the punishers once this minimum number is reached. The sanctioning institution must be maintained regardless of the presence or absence of offenders Perc_2012 . In addition, and in contrast to most studies of public goods games that consider small groups of players, we focus on large groups (i.e., communities) for which sanctioning institutions are pervasive.
In the infinite population limit, where the game can be studied using the replicator equation framework Hofbauer_1998 , we find that the all-defectors solution is always stable and the all-cooperators solution is always unstable. An equilibrium solution where the population consists only of institutional punishers is stable if the per capita cost to punishers of maintaining the sanctioning institution (the maximum per capita cost is $\gamma$) is less than the fine to cooperators ($\alpha\beta$) and this cost plus the cost $c$ of contributing to public goods is less than the fine to defectors ($\beta$). In this bistability scenario, there is a lower value of the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders (i.e., cooperators and defectors). This threshold has a non-trivial dependence on the size of the community. These findings are consistent with the results of other implementations of institutional punishment where both cooperators and defectors are punished Szolnoki_2011 ; Perc_2012 ; Dercole_2013 .
For finite populations, the game is studied using the imitation stochastic dynamics, which reproduces the results of the replicator equation framework in the infinite population limit Fontanari_2024b ; Traulsen_2005 ; Sandholm_2010 . In the case the fine to cooperators is too small, so they win over the punishers in the deterministic regime, we use finite-size scaling Binder_1985 ; Privman_1990 ; Campos_1999 to show that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size.
The rest of the paper is organized as follows. In Section II, we present a variant of the $N$-person prisoner’s dilemma that more realistically models both the maintenance and the operation of sanctioning institutions. In particular, in this section we present the instantaneous payoffs of the three strategies that players can adopt, namely cooperators, defectors, and institutional punishers, and describe the stochastic imitation dynamics that govern the evolution of strategy frequencies for finite populations. In Section III we study the game for infinitely large populations using the replicator equations framework: we find the equilibrium solutions and the conditions for local stability, and explore the dynamics by numerically solving the replicator equations and representing the orbits in the simplex. In Section IV we study the finite population version of the game, using Monte Carlo simulations to implement the stochastic imitation dynamics and finite-size scaling techniques to infer fixation probabilities and mean time to fixation in the limit of very large populations. Finally, in Section V we summarize our results and present some concluding remarks.
II The model
Consider a well-mixed population of finite size $M$, composed of $X$ cooperators, $Y$ defectors, and $Z$ institutional punishers, such that $X+Y+Z=M$. Since we are only considering institutional punishers here, we will refer to these individuals simply as punishers. The population is well mixed in the sense that every individual can interact with every other individual in the population. At each time step $\delta t$, a focal individual $i$ is randomly selected. The play group of $i$ is formed by randomly selecting $N-1\geq 1$ other individuals without replacement from the remaining $M-1$ individuals in the population.
The focal individual’s payoff depends on her own strategy as well as the strategies of the other members of her group. Let us consider each possibility separately.
  1. (a)
    Assume that the focal individual $i$ is a cooperator and her group consists of $I$ other cooperators, $J$ defectors, and $K$ punishers. Here $I,J,K=0,\ldots,N-1$ such that $I+J+K=N-1$. The (instantaneous) payoff of the focal individual is $f_{i}=F_{C}(I+1,J,K)$, with
    $F_{C}(I+1,J,K)$ $=$ $\frac{(I+K)rc}{N-1}-c$ (1)
    $-\alpha\beta\Theta(K-K_{m})$
    where the Heaviside function is $\Theta(x)=1$ if $x\geq 0$ and 00 otherwise. Here we consider the so-called others-only scenario, in which the amplified contribution $rc$ is shared by all other participants in the play group, but not by the contributor Hannelore_2010 (see also Hamilton_1975 ). This is the reason why the numerator of the first term on the rhs of Eq. (1) is missing one contributor (the total number of contributors is $I+1+K$). Moreover, each amplified contribution is shared only by $N-1$ members of the group, since the contributors do not benefit from their own contributions. Hence the factor $N-1$ in the denominator of this term. The second term on the rhs of Eq. (1) is the focal individual’s contribution to public goods, and the third term is the penalty for not contributing to the sanctioning institution. This penalty is a fraction $\alpha\in[0,1]$ of the penalty applied to defectors (see below) and it is only effective if there are at least $K_{m}<N$ individuals (i.e., punishers) contributing to the sanctioning institution. Once this minimum number of punishers is reached, the penalty term becomes constant, in contrast to previous formulations of institutional punishment where the penalty increases linearly with the number of individuals contributing to the sanctioning institution Hannelore_2010 ; Sigmund_2010 , resulting in overpunishment Dercole_2013 . In fact, the constant punishment term with $K_{m}=1$ was proposed in the original study that showed the non-necessity of overpunishment for the establishment of public goods contributors in the institutional punishment scenario Dercole_2013 .
  2. (b)
    Assume that the focal individual $i$ is a defector and her group consists of $I$ cooperators, $J$ other defectors, and $K$ punishers so that $I+J+K=N-1$ with $I,J,K=0,\ldots,N-1$. The (instantaneous) payoff of the focal individual is $f_{i}=F_{D}(I,J+1,K)$, with
    $F_{D}(I,J+1,K)$ $=$ $\frac{(I+K)rc}{N-1}$ (2)
    $-\beta\Theta(K-K_{m}).$
    The terms on the rhs of this equation have a similar interpretation to that given in the previous item. We just note that the penalty for defectors is a factor $1/\alpha$ higher than the penalty for cooperators when $K\geq K_{m}$.
  3. (c)
    Assume that the focal individual $i$ is a punisher and her group consists of $I$ cooperators, $J$ defectors, and $K$ other punishers. As before, $I,J,K=0,\ldots,N-1$ such that $I+J+K=N-1$. The (instantaneous) payoff of the focal individual is $f_{i}=F_{P}(I,J,K+1)$, with
    $F_{P}(I,J,K+1)$ $=$ $\frac{(I+K)rc}{N-1}-c-\gamma$ (3)
    $+\gamma\Theta(K+1-K_{m})$
    $-\gamma\frac{K_{m}}{K+1}\Theta(K+1-K_{m}).$
    Only the last term in the rhs of this equation needs explanation. There is a minimum cost of creating the sanctioning institution, which can only be covered by the contribution of $K_{m}$ punishers. This is similar to the variant of the NPD where a minimum number of cooperators is needed to produce the public goods Pacheco_2009 . Once this threshold cost is reached, the punishers optimize their payoffs by sharing among them the costs of maintaining the sanctioning institution, a feature that makes this variant of the NPD similar to the $N$-person snowdrift game Zheng_2007 ; Santos_2012 ; Fontanari_2024a . Besides the introduction of the threshold $K_{m}$, another important difference between our approach and that of Ref. Dercole_2013 is that punishers always contribute to the sanctioning institution, i.e., they contribute even if there are no defectors and cooperators. In fact, their contribution does not depend on the number of offenders at all. These assumptions are more consistent with the way sanctioning institutions are maintained by contributors in the real world.
    To avoid the paradoxical situation that occurs for small $\gamma$, where the cost of setting up the sanctioning institution is negligible, but $K_{m}$ contributors are still needed to punish offenders, we could choose the per capita cost of maintaining the sanctioning institution when the number of punishers is less than $K_{m}$ to be greater than or equal to the contribution to public goods, i.e., $\gamma\geq c$. Although this is a reasonable and defensible assumption, we will allow $\gamma$ to increase from 00 for a better visualization of our results.
Once the focal individual’s payoff $f_{i}$ is determined, a model individual $j\neq i$ is randomly selected along with her play group, which may include the focal individual $i$ or other members of $i$’s play group. The payoff $f_{j}$ of the model individual is obtained using the same procedure described above. If $f_{j}>f_{i}$, then the focal individual $i$ switches to the strategy of the model individual $j$ with probability
$\frac{f_{j}-f_{i}}{\Delta f_{\max}},$ (4)
where $\Delta f_{\max}$ is the maximum possible payoff difference that guarantees that this ratio is less than or equal to 1. If $f_{j}\leq f_{i}$, then the focal individual $i$ keeps her strategy. In our variant of the NPD, we have
$\Delta f_{\max}=(r+1)c+\gamma,$ (5)
which is the difference between the payoff obtained by a single defector in a group of cooperators (i.e., $rc$) and the payoff obtained by a single punisher in a group of defectors (i.e., $-c-\gamma$). This completes the time step $\delta t$ of the stochastic dynamics, and so the time is updated to $t=t+\delta t$. As usual in such an asynchronous update scheme, we choose the time increment to be $\delta t=1/M$, so that during the increment from $t$ to $t+1$ exactly $M$, though not necessarily different, individuals are selected as focal individuals. The particular choice of switching probability given in Eq. (4) produces the replicator equation in the limit $M\to\infty$ Fontanari_2024b (see Traulsen_2005 ; Sandholm_2010 for other switching probabilities that produce the replicator equation in the large population limit).
III The replicator equation formulation
Here we consider an infinitely large population (i.e., $M\to\infty$) consisting of a fraction $x$ of cooperators, a fraction $y$ of defectors and a fraction $z$ of punishers, with $x+y+z=1$. Only the expected payoffs go into the replicator equation Hofbauer_1998 , so we need to calculate them for each of the three strategies. In particular, the expected payoff of a cooperator is
$\pi^{C}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{C}(I+1,J,K)\delta_{I+J+K,N-1}$ (6)
$=$ $rc(x+z)-c$
$-\alpha\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right]$
and the expected payoff of a defector is
$\pi^{D}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{D}(I,J+1,K)\delta_{I+J+K,N-1}$ (7)
$=$ $rc(x+z)$
$-\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right].$
Calculating the expected payoff of a punisher is a bit more complicated and yields
$\pi^{P}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{P}(I,J,K+1)\delta_{I+J+K,N-1}$ (8)
$=$ $rc(x+z)-c$
$-\frac{\gamma K_{m}}{Nz}\left[1-\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}\right]$
$-\gamma\sum_{K=0}^{K_{m}-2}\binom{N-1}{K}z^{K}(1-z)^{N-1-K},$
where it is implicit that the last term in the rhs is zero if $K_{m}=1$.
Sanctioning institutions only make sense for very large play groups, otherwise peer punishment or small coalitions are more appropriate Ostrom_1990 . It is therefore of interest to obtain the expressions for the payoffs of the strategies in the limits $N\to\infty$ and $K_{m}\to\infty$, such that the ratio
$\zeta=\frac{K_{m}}{N}$ (9)
is nonzero. This can be easily achieved by replacing the binomial distributions by Gaussian distributions in Eqs. (6), (7) and (8) and doing the integrations over the number of punishers. The final result is
$\pi^{C}=\begin{cases}rc(x+z)-c&\text{if $z<\zeta$ }\\ rc(x+z)-c-\alpha\beta&\text{otherwise},\end{cases}$ (10)
$\pi^{D}=\begin{cases}rc(x+z)&\text{if $z<\zeta$ }\\ rc(x+z)-\beta&\text{otherwise},\end{cases}$ (11)
and
$\pi^{P}=\begin{cases}rc(x+z)-c-\gamma&\text{if $z<\zeta$ }\\ rc(x+z)-c-\gamma\zeta/z&\text{otherwise}.\end{cases}$ (12)
The replicator equations governing the evolution of the frequency cooperators, defectors and punishers in the infinite population are
$\frac{dx}{dt}$ $=$ $\frac{x}{\Delta f_{\max}}\left[\pi^{C}(x,z)-\bar{\pi}(x,y,z)\right]$ (13)
$\frac{dy}{dt}$ $=$ $\frac{y}{\Delta f_{\max}}\left[\pi^{D}(x,z)-\bar{\pi}(x,y,z)\right]$ (14)
$\frac{dz}{dt}$ $=$ $\frac{z}{\Delta f_{\max}}\left[\pi^{P}(x,z)-\bar{\pi}(x,y,z)\right],$ (15)
where $\bar{\pi}=x\pi^{C}+y\pi^{D}+z\pi^{P}$ is the average payoff of the population and $\Delta f_{\max}$ is given by Eq. (5). Although we could eliminate the factor $1/\Delta f_{\max}$ by a proper rescaling of $t$ and thus obtain the standard form of the replicator equation, keeping it facilitates comparison with the finite population simulations presented in Section IV. We refer the reader to Ref. Hofbauer_1998 for a thorough introduction to the replicator equation and to Ref. Fontanari_2024b for the derivation of the replicator equation as the deterministic limit of the stochastic imitation dynamics introduced in Section II. Since $x+y+z=1$ and the expected payoffs of the different strategies depend only on $x$ and $z$, henceforth we will eliminate $y$ in favor of these variables. The resulting equations become
$\frac{dx}{dt}$ $=$ $\frac{x}{\Delta f_{\max}}\left[(1-x)(\pi^{C}-\pi^{D})-z(\pi^{P}-\pi^{D})\right]$ (16)
$\frac{dz}{dt}$ $=$ $\frac{z}{\Delta f_{\max}}\left[(1-z)(\pi^{P}-\pi^{D})-x(\pi^{C}-\pi^{D})\right]$ (17)
where we have omitted the dependence of the expected payoffs of the strategies on $x$ and $z$ for simplicity. Since only payoff differences appear in these equations, we can drop the common term $rc(x+z)$ in Eqs. (6), (7) and (8), which greatly facilitates our analysis of the equilibrium solutions. On the one hand, as $z$ increases from 00 to $1$, $\tilde{\pi}^{C}(z)=\pi^{C}(x,z)-rc(x+z)$ monotonically decreases from $-c$ to $-c-\alpha\beta$. Similarly, $\tilde{\pi}^{D}(z)=\pi^{D}(x,z)-rc(x+z)$ monotonically decreases from 00 to $-\beta$. On the other hand, $\tilde{\pi}^{P}(z)=\pi^{P}(x,z)-rc(x+z)$ monotonically increases from $-c-\gamma$ to $-c-\gamma\zeta$, as $z$ increases from 00 to $1$. Note that the introduction of these shifted payoffs makes it clear that the equilibrium strategy frequencies do not depend on the amplification factor $r$. In addition, by properly rescaling $t$ we can set $c=1$ in the numerical analysis without loss of generality.
The fixed-point solutions of the replicator equations (16) and (17) are obtained by setting $dx/dt=dz/dt=0$. In the following we briefly present these solutions together with the results of a standard local stability analysis Britton_2003 ; Murray_2007 .
III.1 All-cooperators solution
Clearly, $x=1$ and consequently $y=z=0$ is a solution of $dx/dt=dz/dt=0$ and corresponds to a population consisting only of cooperators. This solution is locally stable provided that the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the $2\times 2$ community matrix are negative. They are
$\lambda_{1}$ $=$ $c$ (18)
$\lambda_{2}$ $=$ $-\gamma.$ (19)
Since $\lambda_{1}>0$ and $\lambda_{2}<0$, the all-cooperators solution is a saddle point. This means that a population of cooperators can be invaded by defectors, but not by punishers. The population mean payoff $\bar{\pi}=c(r-1)$ is the largest possible payoff for a population at equilibrium.
III.2 All-defectors solution
In this case $y=1$ and consequently $x=z=0$. This is clearly a solution of $dx/dt=dz/dt=0$. The eigenvalues of the corresponding community matrix are
$\lambda_{1}$ $=$ $-c$ (20)
$\lambda_{2}$ $=$ $-c-\gamma,$ (21)
which are always negative and so a population of defectors cannot be invaded, i.e., the all-defectors solution is an evolutionarily stable equilibrium Maynard_1982 . However, the population mean payoff $\bar{\pi}=0$ is the worst possible payoff for a population in equilibrium, which reveals the origin of the public goods dilemma: although the highest individual payoff is obtained by a defector in a group of cooperators ($rc$), if every player decides to defect, the population as a whole ends up in the worst possible situation.
III.3 All-punishers solution
In this case $z=1$ and consequently $x=y=0$. As in the previous cases, this is clearly a solution of $dx/dt=dz/dt=0$ and the eigenvalues of the corresponding community matrix are
$\lambda_{1}$ $=$ $-\alpha\beta+\gamma\zeta$ (22)
$\lambda_{2}$ $=$ $\gamma\zeta+c-\beta.$ (23)
Now things get more interesting. The stability condition $\lambda_{1}<0$ means that the per capita cost to punishers of contributing to the sanctioning institution must be less than the penalty for cooperators, while $\lambda_{2}<0$ means that the cost of contributing to public goods and to the sanctioning institution must be less than the penalty for defectors. These are reasonable conditions that are, in principle, likely to be satisfied in real societies that use sanctioning institutions to deter offenders.
III.4 No-cooperators solution
The scenario with no cooperators, i.e. $x=0$ and $y=1-z^{*}$, is a solution of $dx/dt=dz/dt=0$, provided that $z^{*}\in(0,1)$ is a root of $g(z)=\tilde{\pi}^{D}(z)-\tilde{\pi}^{P}(z)$. Since $g(z)$ decreases monotonically from $c+\gamma$ to $-\beta+c+\gamma\zeta$ as $z$ increases from 00 to $1$, it is clear that the single valid root exists only if $-\beta+c+\gamma\zeta<0$, which is the only condition for the stability of the all-punishers fixed point in the absence of cooperators. Therefore, the no-cooperators solution is unstable and the value of $z^{*}$ delimits the basins of attraction of the all-punishers ($z=1$) and the all-defectors ($y=1$) solutions. More precisely, if $z(0)>z^{*}$ the dynamics is driven to the all-punishers fixed point and to the all-defectors fixed point if $z(0)<z^{*}$. So the larger $z^{*}$ is, the smaller is the size of the domain of attraction of the all-punishers solution. Figure 1 shows $z^{*}$ as function of $\gamma$ for different values of the ratio $\zeta$. As expected, the domain of attraction of the all-punishers fixed point decreases with increasing $\gamma$ and $\zeta$. Note that $z^{*}$ does not depend on $\alpha$ since the cooperators play no role in this scenario.
Figure 1: Unstable fixed point $z^{*}$ corresponding to the no-cooperators solution as a function of the maximum per capita contribution to the sanctioning institution $\gamma$. Left panel: $N=50$ and (from bottom to top) $\zeta=0.1,0.2,0.4$ and $0.8$. Right panel: $\zeta=0.2$ and (from top to bottom at $\gamma=2$) $N=50,100,200$ and $400$. The thick curves are the results for finite $N$, while the thin lines are the results for $N\to\infty$, which exist only for $z^{*}\geq\zeta$. The other parameters are $\beta=3$ and $c=1$.
While for finite $N$ we have to obtain the root $z^{*}$ of $g(z)$ numerically, for $N\to\infty$ it can be obtained easily using Eqs. (11) and (12), leading to
$z^{*}=\frac{\gamma\zeta}{\beta-c}$ (24)
provided that $\gamma\geq\beta-c$, which ensures that $z^{*}\geq\zeta$ and so that the offenders are punished. For $\gamma<\beta-c$ there is no root for $g(z)$, which means that the size of the domain of attraction of the all-punishers solution vanishes. As expected, Fig. 1 shows the strong finite $N$ effects near the discontinuity predicted by the analytical solution in the limit $N\to\infty$ as well as for $\gamma<\beta-c$.
III.5 No-defectors solution
The analysis of the scenario without defectors is very similar to that presented above. In particular, $y=0$ and $x=1-z^{*}$ is a solution of $dx/dt=dz/dt=0$, provided that $z^{*}\in(0,1)$ is a root of $h(z)=\tilde{\pi}^{C}(z)-\tilde{\pi}^{P}(z)$. Using the same reasoning as before, we conclude that the single valid root exists only if $-\alpha\beta+\gamma\zeta<0$, which is the only condition for the stability of the all-punishers fixed point in the absence of defectors. We recall that in this case the all-cooperators fixed point is stable. Therefore, the no-defectors solution is unstable and the value of $z^{*}$ delimits the basins of attraction of the all-punishers ($z=1$) and the all-cooperators ($x=1$) solutions. As before, $z^{*}$ can be obtained easily using Eqs. (10) and (12) for $N\to\infty$, leading to
$z^{*}=\frac{\gamma\zeta}{\alpha\beta}$ (25)
provided that $\gamma\geq\beta\alpha$ which ensures that $z^{*}\geq\zeta$. Of course, the results for both the finite and infinite play group sizes depend on the product $\alpha\beta$ and not on the individual values of the parameters $\alpha$ and $\beta$. In addition, the parameter $c$ plays no role in the determination of $z^{*}$ for the no-defectors solution. Since the dependence of $z^{*}$ on $\gamma$ is very similar to that shown in Fig. 1, here we look at how the finite play group solutions approach the discontinuous result predicted by the analytical solution in the limit $N\to\infty$. The left panel of Fig.  2 shows the unstable fixed point $z^{*}$ for fixed $\zeta=0.2$ and $\alpha\beta=4$, and a variety of large play group sizes $N$. The right panel shows that the deviation from the $N\to\infty$ prediction vanishes like $1/\sqrt{N}$ as $N$ increases. This is expected since for large $N$ we have used the approximation
$\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}$$\approx$ $\frac{1}{2}\mbox{erf}\left[\frac{(\zeta-z)\sqrt{N}}{\sqrt{2z(1-z)}}\right]+$ (26)
$\frac{1}{2}\mbox{erf}\left[\frac{z\sqrt{N}}{\sqrt{2z(1-z)}}\right]$
to derive Eqs. (24) and (25). We note that finding how the deviations scale with $N$ by ‘collapsing’ the curves for different values of $N$ into a single curve is a technique widely used in finite-size scaling analysis to calculate the critical exponents of phase transitions Binder_1985 ; Privman_1990 ; Campos_1999 .
Figure 2: The no-defectors unstable fixed point $z^{*}$ as a function of the maximum per capita contribution to the sanctioning institution $\gamma$ for $\zeta=0.2$, $\alpha\beta=4$ and (from top to bottom at $\gamma=3$) $N=3200,6400,12800$ and $25600$ (left panel). The black thin line is the result for $N\to\infty$. The right panel shows the same curves with the y-axis rescaled to $(z^{*}-\zeta)N^{1/2}$.
Figure 3: Values of $\alpha$ and $\gamma$ (left panel) for which the coexistence fixed point $z^{*}$ (right panel) exists for $\zeta=0.2$, and (from top to bottom at $\alpha=0.6$) $N=50,200,1600$ and $3200$. The other parameters are $\beta=3$ and $c=1$. The vertical dashed lines indicate the threshold $\alpha=1-c/\beta=2/3$ beyond which there is no coexistence solution. The black thin curves indicate the results for $N\to\infty$.
III.6 Coexistence solution
The existence of an equilibrium solution where all three strategies are present would require a value of $z^{*}$ such that $g(z^{*})=\tilde{\pi}^{D}(z^{*})-\tilde{\pi}^{P}(z^{*})=0$ and $h(z^{*})=\tilde{\pi}^{C}(z^{*})-\tilde{\pi}^{P}(z^{*})=0$. Solving these equations simultaneously, we find that such values do indeed exist for certain values of the model parameters, as shown in Fig. 3. However, the above equilibrium conditions only determine the value of $z^{*}$, so $x^{*}$ must depend on the initial conditions. In fact, if we set $z(0)=z^{*}$, then the replicator dynamics freezes at $x^{*}=x(0)$, which means that there are infinitely many coexistence fixed points.
The derivation of the results for $N\to\infty$ is a bit more complicated than in the previous cases. The reason is that as $N$ increases, $z^{*}$ approaches $\zeta$ so that $(\zeta-z^{*})N^{1/2}=\eta$ is finite and nonzero. Using the approximation (26) we can write the equations $g(z^{*})=h(z^{*})=0$ as
$\alpha=1-\frac{c}{\beta[1-H(\eta)]}$ (27)
$\gamma=\frac{c\alpha}{1-\alpha}$ (28)
where
$H(\eta)=\frac{1}{2}+\frac{1}{2}\mbox{erf}\left[\frac{\eta}{\sqrt{2\zeta(1-\zeta)}}\right].$ (29)
For fixed $\beta$, the largest value of $\alpha$ for which Eq. (27) admits a solution is $\alpha=1-c/\beta$, corresponding to $\eta\to-\infty$, and in this case we have $\gamma=\beta-c$. The left panel of Fig. 3 shows the very good agreement between Eq. ( 28) and the numerical results for large $N$. Of course, whenever Eq. (27) admits a solution we have $z^{*}=\zeta$, which is indeed the large $N$ limit of the numerical results, as shown in the right panel of Fig. 3.
This completes the analysis of the fixed points of the replicator equations (16) and (17). Note that there is no fixed point corresponding to the no-punishers scenario (i.e., $z=0$ and $y=1-x$ with $x\in(0,1)$) since $\tilde{\pi}^{D}(0)>\tilde{\pi}^{C}(0)$ for all parameter settings.
III.7 Dynamics
The all-defectors fixed point ($y=1$) is stable for all values of the model parameters, while the all-punishers fixed point ($z=1$) is stable if the conditions $\gamma\zeta<\alpha\beta$ and $\gamma\zeta<\beta-c$ are satisfied. Except for the coexistence and for the no-cooperators fixed points, which are unstable (see figures below), all other fixed points and especially the all-cooperators fixed point ($x=1$) are saddle points. Thus, the useful information we can obtain from the study of the replicator dynamics, i.e., the numerical solution of Eqs. (16) and (17), is the domains of attraction of the stable fixed points, which are better visualized by ternary or simplex plots. In these plots, the vertices of the triangle denote homogeneous populations of punishers (P), cooperators (C), and defectors (D).
Figure 4 shows the replicator dynamics in the case of bistability of the all-punishers and all-defectors fixed points. The starting points of the orbits are very close to the unstable fixed points. The left panel illustrates the scenario where the coexistence fixed points exist (the model parameters must be set very precisely according to Fig. 3): since the coexistence condition only fixes the frequency of punishers $z^{*}$, there are infinitely many unstable coexistence fixed points determined by the different values of the initial frequency of cooperators $x^{*}=x(0)$. Of course, the line of the unstable fixed point coincides with the separatrix delimiting the domains of attraction of the all-punishers and all-defectors fixed points. The right panel illustrates the more common bistability scenario, where there are no coexistence fixed points and the boundary of the attraction domains is determined by a separatrix, which directs the flow toward the no-defectors fixed point. Since the only difference between the two panels is the value of $\alpha$, the no-cooperators fixed point is the same in both plots. In fact, we can now see that this fixed point is unstable, whereas the no-defectors fixed point is a saddle. This is only true for the scenario of bistability between the all-punishers and the all-defectors fixed point, otherwise the no-defectors fixed point is also unstable.
Figure 4: Replicator dynamics for (left panel) $\alpha=0.549$ and $\gamma=1.499$ and (right panel) $\alpha=0.2$ and $\gamma=1.499$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the fixed points. The blue orbits converge to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbit converges to the all-cooperators fixed point. The magenta horizontal line in the left panel is a line of (unstable) coexistence fixed points and is the separatrix delimiting the domains of attraction of P and D. The magenta curve in the right panel is the separatrix over which the dynamics goes to the no-defectors fixed point. The black dashed line is $z=\zeta=0.2$.
Figure 5 shows the scenarios where the all-punisher fixed point is a saddle and thus the all-defectors fixed point is the only stable solution of the replicator dynamics. The left panel illustrates the case $\gamma\zeta>\alpha\beta$ and $\gamma\zeta<\beta-c$ so that the cooperators win over the punishers. A typical orbit leads to an almost complete dominance of the population by cooperators, which are then quickly overrun by defectors since the small number of punishers is not enough to establish a sanctioning institution. The right panel illustrates the case $\gamma\zeta<\alpha\beta$ and $\gamma\zeta>\beta-c$ so that the defectors win over the punishers. As before, the starting points of the orbits are close to the unstable fixed points at the edges of the simplex. The frequency of both cooperators and defectors increases at the expense of the punishers, but once they disappear, the defectors quickly take over the population.
Figure 5: Replicator dynamics for (left panel) $\alpha=0.05$ and $\gamma=1.5$ and (right panel) $\alpha=0.9$ and $\gamma=12.5$. The other parameters are $N=50$, $K_{m}=10$, $\beta=3$ and $c=1$. The filled circles indicate the fixed points. The blue orbit converges to the all-punishers fixed point, the red orbits converge to the all-defectors fixed point, and the green orbits converge to the all-cooperators fixed point. The black dashed line is $z=\zeta=0.2$.
The regime of bistability between the all-punishers and the all-defectors fixed points is the more interesting feature of the replicator dynamics. The resulting scenario can be better appreciated by the separatrix curves that delimit the domains of attraction of these fixed points. These curves are shown in Fig. 6 for a variety of values of the parameters $\alpha$ (left panel) and $\beta$ (right panel) such that $\gamma\zeta<\alpha\beta$ and $\gamma\zeta<\beta-c$. Note that $z^{*}=\zeta$ is a lower bound on the frequency of punishers only in the limit $N\to\infty$. For finite $N$, there is a nonzero probability of play groups with $K>K_{m}$ punishers even if $z<\zeta$ (see Figs. 1, 2, and 3).
Figure 6: Separatrices delimiting the domains of attraction of P and D for $\gamma=1.5$ and (left panel from top to bottom) $\alpha=0.11,012,0.15,0.2,0.3,1$ and $\beta=3$, and (right panel from top to bottom) $\beta=1.6,2,3,4,5,10$ and $\alpha=0.2$. The other parameters are $N=50$, $K_{m}=10$, and $c=1$. The filled circles indicate the fixed points. The black dashed line is $z=\zeta=0.2$. The flow over a separatrix is always towards the no-defectors fixed point.
Figures 4 and 5 illustrate the typical dynamic scenarios that hold for finite and infinite play group sizes. In particular, the right panel of Fig. 4 makes clear that in the presence of cooperators a larger number of punishers is required to eliminate the threat of defectors, since the no-defectors fixed point occurs at a larger value of z than the no-cooperators fixed point. The left panel of Fig. 5 shows the detrimental effect of the presence of improperly punished cooperators: they exploit the rarity of defectors to overrun the punishers, opening the way for the invasion of defectors.
IV Finite population simulations
As already pointed out, a major advance in the theoretical study of public goods games is the realization that the replicator equation formalism originally introduced to study biological evolution in continuous time Hofbauer_1998 also describes the social dynamics scenario where individuals are more likely to imitate the behavior of their more successful peers Fontanari_2024b . In this section, we offer a comparison between the predictions of the deterministic replicator equations (16) and (17) and the finite population simulations using the stochastic imitation dynamics presented in Section II.
Although the probability that the focal individual switches strategy, Eq.  (4), depends on the difference between her instantaneous payoff and the payoff of the model individual, their shares of the public goods do not cancel out because they are evaluated for different play groups. Thus, in principle, the simulation results could depend on the amplification factor $r$. Of course, we expect this effect to be negligible for large population sizes $M$.
Figure 7: Frequencies of cooperators $x$, defectors $y$ and punishers $z$ as a function of time for the initial conditions $z(0)=0.25$ and $x(0)=0.1$ (left panel), $x(0)=0.11052$ (middle panel) and $x(0)=0.3$ ( right panel). The thin curves are the four runs of the stochastic simulation algorithm for $M=10^{4}$ and the thick curves are the numerical solutions of the replicator equations. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.2$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
Figure 7 shows the numerical solution of Eqs. (16) and (17) together with four independent runs of the stochastic simulation algorithm for populations of size $M=10^{4}$ and the same parameters of the simplex plot in the right panel of Fig. 4. The finite-size effects in the dynamics are very small for initial conditions far from the separatrix (left and right panels). These two panels illustrate the nonintuitive result that increasing the number of cooperators while keeping the number of punishers fixed (and thus decreasing the number of defectors) leads to the dominance of defectors. The middle panel shows the dynamics when the initial conditions $x(0)$ and $z(0)$ are very close to the separatrix. If it were possible to choose these initial values exactly at the separatrix and to evolve the deterministic dynamics without round-off errors, then the dynamics should lead to the no-defectors fixed point, $x^{*}=z^{*}=0.5$ for the parameter setting of the figure. In fact, if these small numerical errors eventually move the deterministic trajectory away from the no-defectors fixed point, then we expect the effect of the intrinsic noise of the finite population simulations to be much more disruptive. This is indeed the case: among the four runs of the stochastic algorithm, two lead to the fixation of the punishers, one to the fixation of the cooperators, and one to the fixation of the defectors. Most interestingly, since the orbits leading to the all-defectors fixed point may be too close to the no-defectors edge of the simplex (see the right panel of Fig. 4 and the left panel of Fig. 5), the demographic noise is likely to favor the fixation of the cooperators.
Figure 8: Probability of fixation of cooperators $\Pi_{C}$ as a function of their initial frequency (left panel) for populations of sizes (from top to bottom) $M=100,500,1000,5000,10000$ and $20000$. The right panel shows $\Pi_{C}$ as a function of the scaled variable $x(0)M^{0.15}$ for $M=5000,10000$ and $20000$. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
To quantify the advantage of the cooperators for finite populations, we consider the parameter setting of the simplex plot in the left panel of Fig. 5, where any orbit starting from an initial condition inside the simplex leads to the all-defectors fixed point. The probabilities of fixation $\Pi_{C}$, $\Pi_{D}$, and $\Pi_{P}$ are approximated by the fractions of the runs that lead to the fixation of cooperators, defectors and punishers, respectively. In Fig.  8 we fix $z(0)=0.5$ and vary $x(0)$ in the interval $(0,0.5)$ for several population sizes. The total number of independent runs is $10^{4}$. Since we never observed the fixation of the punishers (i.e., $\Pi_{P}=0$), it is sufficient to show $\Pi_{C}$ in this figure. The results show that fixation of cooperators is very likely for small populations. Even for large populations, the cooperators are very likely to end up as winners, provided, somewhat counterintuitively, that their initial frequency is sufficiently low. This is a consequence of the peculiarity of the trajectories shown in the simplex plot in the left panel of Fig. 5. As $M$ increases, the range of initial conditions $x(0)$ for which $\Pi_{C}$ is not zero shrinks to zero, as expected from the analysis of the replicator equations. More specifically, $x(0)$ decreases with $M^{-0.15}$, as shown in the right panel of Fig. 8. This result implies that, for example, for $x(0)=0.2M^{-0.15}$ and $z(0)=0.5$, the cooperators are virtually certain to fixate regardless of the size of the population. Another important quantity to characterize the stochastic dynamics is the mean fixation time $T_{f}$, i.e., the mean time for the dynamics to reach the all-cooperators or the all-defectors absorbing configurations, which is shown in Fig. 9. For small $x(0)$ and large $M$, the cooperators fixate in practically all runs (see Fig. 8) and the fixation time decreases with increasing $x(0)$, but undergoes an abrupt increase when it is the turn of the defectors to fixate with certainty. In this regime, $T_{f}$ decreases slowly as $x(0)$ increases, which implies that $y(0)$ decreases. In the region of interest, i.e., for $\Pi_{C}\approx 1$ and large $M$, the scaling assumption
$T_{f}=M^{b}f\left[x(0)M^{a}\right]$ (30)
with $a=0.15$ and $b=0.1$ fits the data very well, as shown in the right panel of Fig. 9. Here $f(u)$ is a scaling function that satisfies $\lim_{u\to 0}f(u)\to\infty$. The low values of the exponents $a$ and $b$ in this scaling relation indicate the small quantitative effect of the population size on $\Pi_{C}$ and $T_{f}$.
Figure 9: Mean time for fixation of cooperators or defectors $T_{f}$ as a function of the initial frequency of cooperators (left panel) for populations of sizes (from bottom to top) $M=100,500,1000,5000,10000$ and $20000$. The right panel shows the suitability of the scaling assumption (30) for $M=5000,10000$ and $20000$. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
To understand how the fixation probability of the cooperators is affected by the choice of the initial frequency of punishers, in Fig. 10 we set $x(0)=0.2M^{-0.15}$ and vary $z(0)$ for different population sizes $M$. As shown before, this scaling guarantees the fixation of the cooperators for $z(0)=0.5$ in the limit of large $M$. The results indicate that there is a threshold for $z(0)$ below which the cooperators have no chance against the defectors. At this threshold, the mean fixation time is maximal. The finite size effects on these properly scaled measures are negligible, except very close to the threshold.
Figure 10: Probability of fixation of cooperators $\Pi_{C}$ (left panel) and scaled mean time for fixation $T_{f}/M^{0.1}$ (right panel) as a function of initial frequency of punishers $z(0)$ for $M=5000,10000$ and $20000$. The initial frequency of cooperators is set to $x(0)=0.2M^{-0.15}$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
The finite population analysis gives unexpected results only for the scenario shown in the left panel of Fig. 5, where the replicator equations predict that the cooperators win over the punishers. In the bistability scenario of the right panel of Fig. 4, the initial conditions must be set very close to the separatrix in order for the results of the stochastic dynamics to differ from the deterministic predictions, as shown in Fig. 7. As expected, in the scenario where the defectors win over the punishers in the deterministic regime, shown in the right panel of Fig. 5, the result of the stochastic dynamics is the fixation of the defectors.
V Discussion
In contrast to most public goods game studies, which, following the seminal experimental studies in behavioral economics Fehr_2000 ; Fehr_2002 , consider small groups of players, we focus here on large groups (i.e., communities), since the setting up and maintenance of a sanctioning institution requires a large number of committed individuals. Peer punishment (i.e., vigilantism) is unlikely in communities not only because of the high individual cost of exposure to retaliation, but also because of a well-known phenomenon in social psychology known as the diffusion of responsibility effect: people often fail to cooperate or punish offenders because they believe that others will or should take responsibility for doing so Darley_1968 .
In the community context, we note the striking similarity between the replicator equation approach to public goods games and Wilson’s model of natural selection of populations and communities Wilson_1980 , which focuses on individual traits that affect the fitness of other members of the population, such as behaviors that alter the environment (e.g., pollution and resource depletion). More pointedly, in Wilson’s model the fitness of individuals is determined locally, taking into account interactions within their trait groups, but their chances of reproduction are dictated by competition in the population at large. This is virtually identical to the imitation dynamics of the $N$-person games: a player’s payoff is determined by the composition of the group of $N$ players to which she belongs, but the decision to change her strategy or not is determined by comparison with the payoff of another player randomly selected from the population, as described in Section II. Another point of similarity between the more recent interdisciplinary game theoretic approach to social dilemmas and the traditional biological approach to the evolution of altruism that is worth noting is that the payoffs of cooperators and defectors in the others-only scenario in the absence of punishment (i. e, setting $\beta=0$ in Eqs. (1) and (2)) are identical to the fitness used in Hamilton’s classic work on the innate social behavior of humans Hamilton_1975 , a point overlooked in the recent literature Hannelore_2010 .
In addition to the emphasis on communities, we consider some realistic improvements to the standard formulation of the $N$-person prisoner’s dilemma with punishment, namely the elimination of overpunishment (i.e., the penalties for the different types of offenses are fixed), the high cost of sanctioning institutions (hence a minimum number $K_{m}=\zeta N$ of punishers to set them up), which must be maintained regardless of the presence or absence of offenders, and the division of the maintenance costs among the punishers.
As expected, the replicator equations which describe the imitation dynamics in the infinite population limit (i.e., $M\to\infty$) predict that the all-defectors solution is always stable and that the all-cooperators solution is always unstable. The all-punishers solution is stable if the per capita cost to punishers of contributing to the sanctioning institution is less than the penalty to cooperators, and this cost plus the cost of contributing to public goods is less than the penalty to defectors. In this case, there is a lower threshold $z^{*}$ for the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders, i.e., cooperators and defectors. This threshold has a non-trivial dependence on the community size $N$ (see Figs. 1 and 2): $z^{*}$ increases with increasing $N$ when the per capita contribution to the sanctioning institution $\gamma$ is small, and decreases when it is large. For infinitely large communities (i.e., $N\to\infty$), we find $z^{*}\geq\zeta$, as expected.
The imitation stochastic dynamics Fontanari_2024b ; Traulsen_2005 ; Sandholm_2010 for finite populations produces some unexpected results in the case where the replicator equations predict that cooperators win over punishers, who in turn win over defectors if their initial frequency is sufficiently high (see left panel of Fig 5). We find that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size. Of course, since a population of cooperators can be invaded by a single defector regardless of its size, this is not a robust scenario for maintaining cooperation.
Our approach follows the Hobbesian view that, human beings being the selfish creatures that they are, life would be "solitary, poor, nasty, brutish, and short", were it not for the existence of repressive mechanisms Hobbes_1651 . From a game theory perspective, the main difficulty in establishing this scenario, which seems to be the current state of affairs, is the need for a minimum number of individuals willing to contribute to the establishment of a sanctioning institution. Possible evolutionary trajectories leading to the current repressive dominant system include temporary ritual ceremonies ubiquitous in early human societies, as well as the tendency of human groups to differentiate themselves from neighbors Graeber_2021 . It would be most instructive to incorporate these elements into a game-theoretic framework.
Acknowledgements.
JFF is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico grant number 305620/2021-5.